### v1

@@ 4 Concluding remarks
This paper introduces a practical way to address the parallel trends assumption in DDD models by transforming them into a DID framework and then combining them with synthetic control, in order to improve counterfactual estimation when parallel trends are likely violated across multiple dimensions. This approach is especially relevant in evaluating complex policies such as China’s “4+7 Cities” Centralized Drug Procurement pilot program, where treatment effects may differ significantly between subgroups.
The introduction of synthetic control in DDD settings addresses a key gap in the literature. While balancing methods have been widely applied in DID contexts to account for violations of parallel trends, their application to DDD has been largely unexplored. For empirical researchers, this method expands the use of synthetic control to capture subgroup-specific variations in treatment effects, which are often missed by traditional DDD estimators, and allows for more precise estimation of treatment effects in complex empirical settings where interactions between subgroups and treatment are likely to influence the outcome. By offering a more flexible and robust framework, this approach encourages researchers to move beyond the standard DDD assumptions when subgroup-specific violations of parallel trends are suspected.
@@
While this paper demonstrates the efficacy of combining synthetic control with DDD, there are opportunities for future research. Further work could explore its application in other policy contexts or investigate alternative weighting methods to enhance the performance of synthetic control in DDD settings with complex data structures. This combination of synthetic control and DDD offers a powerful tool for empirical research, expanding the scope of more reliable policy evaluation.
@@ A.1 Transforming the ATT in a simple DID
$\begin{split}\alpha_{1}=&\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=1]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=0]\right)-\\ &\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=1]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=0]\right)\\ =&\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=1]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=1]\right)-\\ &\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=0]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=0]\right)\\ =&\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=1]\mid\text{Treat}_{i}=1,\text{Post}_{t}=1\right]-\\ &\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=0]\mid\text{Treat}_{i}=1,\text{Post}_{t}=0\right]\\ =&\mathbb{E}\left[\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,t=t]\mid\text{Post}_{t}=1\right]-\right.\\ &\left.\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,t=t]\mid\text{Post}_{t}=0\right]\mid\text{Treat}_{i}=1\right]\\ =&\mathbb{E}\left[\mathbb{E}\left[Z_{it}\mid\text{Post}_{t}=1\right]-\mathbb{E}\left[Z_{it}\mid\text{Post}_{t}=0\right]\mid\text{Treat}_{i}=1\right]\end{split}$ (A.1)
@@ A.2 Transforming the triple difference effect
$\begin{split}\delta=&\left[(\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=1,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=0,\text{G}_{i}=1])\right.\\ &-\left.(\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=1,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=0,\text{G}_{i}=1])\right]\\ &-\left[(\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=1,\text{G}_{i}=0]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=0,\text{G}_{i}=0])\right.\\ &-\left.(\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=1,\text{G}_{i}=0]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=0,\text{G}_{i}=0])\right]\\ =&\left[(\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=1,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=1,\text{G}_{i}=0])\right.\\ &-\left.(\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=0,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=0,\text{G}_{i}=0])\right]\\ &-\left[(\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=1,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=1,\text{G}_{i}=0])\right.\\ &-\left.(\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=0,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=0,\text{G}_{i}=0])\right]\\ =&\left[\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=1,\text{P}_{i}=1,\text{G}_{i}=1\right]\right.\\ &-\left.\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=1,\text{P}_{i}=0,\text{G}_{i}=1\right]\right]\\ &-\left[\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=0,\text{P}_{i}=1,\text{G}_{i}=1\right]\right.\\ &-\left.\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=0,\text{P}_{i}=0,\text{G}_{i}=1\right]\right]\\ =&\mathbb{E}\left[\left[\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=1,\text{P}_{i}=1\right]\right.\right.\\ &-\left.\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=1,\text{P}_{i}=0\right]\right]\\ &-\left[\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=0,\text{P}_{i}=1\right]\right.\\ &-\left.\left.\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=0,\text{P}_{i}=0\right]\right]\mid\text{G}_{i}=1\right]\\ =&\mathbb{E}\left[\left[\mathbb{E}\left[W_{it}\mid\text{T}_{i}=1,\text{P}_{i}=1\right]-\mathbb{E}\left[W_{it}\mid\text{T}_{i}=1,\text{P}_{i}=0\right]\right]\right.\\ &-\left.\left[\mathbb{E}\left[W_{it}\mid\text{T}_{i}=0,\text{P}_{i}=1\right]-\mathbb{E}\left[W_{it}\mid\text{T}_{i}=0,\text{P}_{i}=0\right]\right]\mid\text{G}_{i}=1\right]\end{split}$ (A.2)
where $W_{it}$ is defined by Equation (12), T stands for Treat, and P stands for Post.
@@
$W_{it}=\theta_{0}+\theta_{1}\text{D}_{t}+\mu_{i}+\lambda_{t}+e_{it}.$ (A.3)

### v2

@@ 5 Concluding Remarks
This paper introduces a practical way to address the parallel trends assumption in DDD models by transforming them into a DID framework and then combining them with SC, in order to improve counterfactual estimation when parallel trends are likely violated across multiple dimensions. This approach is especially relevant in evaluating complex policies such as China’s “4+7 Cities” Centralized Drug Procurement pilot program, where treatment effects may differ significantly between subgroups.
The introduction of SC in DDD settings addresses a key gap in the literature. While balancing methods have been widely applied in DID contexts to account for violations of parallel trends, their application to DDD has been largely unexplored. For empirical researchers, this method expands the use of SC to capture subgroup-specific variations in treatment effects, which are often missed by traditional DDD estimators, and allows for more precise estimation of treatment effects in complex empirical settings where interactions between subgroups and treatment are likely to influence the outcome. By offering a more flexible and robust framework, this approach encourages researchers to move beyond the standard DDD assumptions when subgroup-specific violations of parallel trends are suspected.
@@
While this paper demonstrates the efficacy of combining SC with DDD, there are opportunities for future research. Further work could explore its application in other policy contexts or investigate alternative weighting methods to enhance the performance of SC in DDD settings with complex data structures. This combination of SC and DDD offers a powerful tool for empirical research, expanding the scope of more reliable policy evaluation.
@@ A.1 Transforming the ATT in a simple DID
$\begin{split}\alpha_{1}=&\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=1]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=0]\right)-\\ &\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=1]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=0]\right)\\ =&\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=1]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=1]\right)-\\ &\left(\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=1,\text{Post}_{t}=0]-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=0]\right)\\ =&\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=1]\mid\text{Treat}_{i}=1,\text{Post}_{t}=1\right]-\\ &\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,\text{Post}_{t}=0]\mid\text{Treat}_{i}=1,\text{Post}_{t}=0\right]\\ =&\mathbb{E}\left[\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,t=t]\mid\text{Post}_{t}=1\right]-\right.\\ &\left.\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{Treat}_{i}=0,t=t]\mid\text{Post}_{t}=0\right]\mid\text{Treat}_{i}=1\right]\\ =&\mathbb{E}\left[\mathbb{E}\left[Z_{it}\mid\text{Post}_{t}=1\right]-\mathbb{E}\left[Z_{it}\mid\text{Post}_{t}=0\right]\mid\text{Treat}_{i}=1\right]\\ =&\mathbb{E}\left[Z_{it}\mid\text{Post}_{t}=1\right]-\mathbb{E}\left[Z_{it}\mid\text{Post}_{t}=0\right]\text{ if }\text{Treat}_{i}=1\end{split}$ (A.1)
@@ A.2 Transforming the triple difference effect
$\begin{split}\delta=&\left[(\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=1,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=0,\text{G}_{i}=1])\right.\\ &-\left.(\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=1,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=0,\text{G}_{i}=1])\right]\\ &-\left[(\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=1,\text{G}_{i}=0]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=0,\text{G}_{i}=0])\right.\\ &-\left.(\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=1,\text{G}_{i}=0]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=0,\text{G}_{i}=0])\right]\\ =&\left[(\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=1,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=1,\text{G}_{i}=0])\right.\\ &-\left.(\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=0,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=1,\text{P}_{t}=0,\text{G}_{i}=0])\right]\\ &-\left[(\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=1,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=1,\text{G}_{i}=0])\right.\\ &-\left.(\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=0,\text{G}_{i}=1]-\mathbb{E}[Y_{it}\mid\text{T}_{i}=0,\text{P}_{t}=0,\text{G}_{i}=0])\right]\\ =&\left[\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=1,\text{P}_{i}=1,\text{G}_{i}=1\right]\right.\\ &-\left.\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=1,\text{P}_{i}=0,\text{G}_{i}=1\right]\right]\\ &-\left[\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=0,\text{P}_{i}=1,\text{G}_{i}=1\right]\right.\\ &-\left.\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=0,\text{P}_{i}=0,\text{G}_{i}=1\right]\right]\\ =&\mathbb{E}\left[\left[\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=1,\text{P}_{i}=1\right]\right.\right.\\ &-\left.\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=1,\text{P}_{i}=0\right]\right]\\ &-\left[\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=0,\text{P}_{i}=1\right]\right.\\ &-\left.\left.\mathbb{E}\left[Y_{it}-\mathbb{E}[Y_{it}\mid\text{T}_{i}=j,t=t,\text{G}_{i}=0]\mid\text{T}_{i}=0,\text{P}_{i}=0\right]\right]\mid\text{G}_{i}=1\right]\\ =&\mathbb{E}\left[\left[\mathbb{E}\left[W_{itj}\mid\text{T}_{i}=1,\text{P}_{i}=1\right]-\mathbb{E}\left[W_{itj}\mid\text{T}_{i}=1,\text{P}_{i}=0\right]\right]\right.\\ &-\left.\left[\mathbb{E}\left[W_{itj}\mid\text{T}_{i}=0,\text{P}_{i}=1\right]-\mathbb{E}\left[W_{itj}\mid\text{T}_{i}=0,\text{P}_{i}=0\right]\right]\mid\text{G}_{i}=1\right]\\ =&\left[\mathbb{E}\left[W_{itj}\mid\text{T}_{i}=1,\text{P}_{i}=1\right]-\mathbb{E}\left[W_{itj}\mid\text{T}_{i}=1,\text{P}_{i}=0\right]\right]\\ &-\left[\mathbb{E}\left[W_{itj}\mid\text{T}_{i}=0,\text{P}_{i}=1\right]-\mathbb{E}\left[W_{itj}\mid\text{T}_{i}=0,\text{P}_{i}=0\right]\right]\text{ if }\text{G}_{i}=1\end{split}$ (A.2)
where $W_{itj}$ is defined by Equation (12), T stands for Treat, and P stands for Post.
@@
$W_{it}=\theta_{0}+\theta_{1}\text{D}_{t}+\mu_{i}+\lambda_{t}+e_{it}\text{ if }G_{i}=1.$ (A.3)
@@ Appendix C More Details about Data
Appendix D An Additional Result
Figure A1: When the parallel trends of synthetic DID are not good enough
Notes: The figure shows how the number of A61 patent applications evolves over time and is affected by the Centralized Drug Procurement pilot program based on a synthetic DID regression with covariates listed in Table 1 using A61 tracks. As shown, the pre-treatment trends are not parallel even after synthetic control.